# From RAG to Memory: Non-Parametric Continual Learning for Large Language Models

### v1

@@ Abstract
Our ability to continuously acquire, organize, and leverage knowledge is a key feature of human intelligence that AI systems must approximate to unlock their full potential. Given the challenges in continual learning with large language models (LLMs), retrieval-augmented generation (RAG) has become the dominant way to introduce new information. However, its reliance on vector retrieval hinders its ability to mimic the dynamic and interconnected nature of human long-term memory. Recent RAG approaches augment vector embeddings with various structures like knowledge graphs to address some of these gaps, namely sense-making and associativity. However, their performance on more basic factual memory tasks drops considerably below standard RAG. We address this unintended deterioration and propose HippoRAG 2, a framework that outperforms standard RAG comprehensively on factual, sense-making, and associative memory tasks. HippoRAG 2 builds upon the Personalized PageRank algorithm used in HippoRAG and enhances it with deeper passage integration and more effective online use of an LLM. This combination pushes this RAG system closer to the effectiveness of human long-term memory, achieving a 7% improvement in associative memory tasks over the state-of-the-art embedding model while also exhibiting superior factual knowledge and sense-making memory capabilities. This work paves the way for non-parametric continual learning for LLMs. Our code and data will be released at https://github.com/OSU-NLP-Group/HippoRAG.
@@ 2.1 Continual Learning for LLMs
Continual learning methods applied to LLMs aim to allow them to acquire and integrate new knowledge over time while preserving past information. Given the high computational cost of full-scale LLM pretraining, various techniques have been used to achieve this goal. These approaches generally fall into three categories: continual fine-tuning, model editing, and RAG (Shi et al., 2024).
@@ 2.2 Non-Parametric Continual Learning for LLMs
Encoder model improvements, particularly with LLM backbones, have significantly enhanced RAG systems by generating high-quality embeddings that better capture semantic relationships, improving retrieval quality for LLM generation. Recent models (Li et al., 2023; Muennighoff et al., 2024; Lee et al., 2025) leverage LLMs, large corpora, improved architectures, and instruction fine-tuning for notable retrieval gains. NV-Embed-v2 (Lee et al., 2025) serves as the primary comparison in this paper.
@@ 3 HippoRAG 2
3.1 Overview
HippoRAG (Gutiérrez et al., 2024) is a neurobiologically inspired long-term memory framework for LLMs, with each component designed to emulate aspects of human memory. The framework consists of three primary components: the artificial neocortex (LLM), the parahippocampal region (PHR encoder), and the artificial hippocampus (open KG). These components collaborate to replicate the interactions observed in human long-term memory.
For HippoRAG offline indexing, an LLM processes passages into KG triples, which are then incorporated into the artificial hippocampal index. Meanwhile, the PHR is responsible for detecting synonymy to interconnect information. For HippoRAG online retrieval, the LLM neocortex extracts named entities from a query, while the PHR encoder link these entities to the hippocampal index. Then, the Personalized PageRank (PPR) algorithm on the KG is conducted for context-based retrieval. Although HippoRAG seeks to construct memory from non-parametric RAG, its effectiveness is hindered by a critical flaw: an entity-centric approach that causes context loss during both indexing and inference, as well as difficulties in semantic matching.
Built on the neurobiologically inspired long-term memory framework proposed in HippoRAG (Gutiérrez et al., 2024), the structure of HippoRAG 2 follows a similar two-stage process: offline indexing and online retrieval, as shown in Figure 2. Additionally, however, HippoRAG 2 introduces several key refinements that improve its alignment with human memory mechanisms: 1) It seamlessly integrates conceptual and contextual information within the open KG, enhancing the comprehensiveness and atomicity of the constructed index (§3.2). 2) It facilitates more context-aware retrieval by leveraging the KG structure beyond isolated KG nodes (§3.3). 3) It incorporates recognition memory to improve seed node selection for graph search (§3.4). In the following sections, we introduce the pipeline in more detail and elaborate on each of these refinements.
Offline Indexing. 1) HippoRAG 2 leverages an LLM to extract triples from each passage and integrates them into a schema-less open KG. We call the subject or object of these triples phrase and the edge connecting them relation edge. 2) Next, the encoder identifies synonyms by evaluating phrase pairs within the KG, detecting those with vector similarity above a predefined threshold, and adding synonym edge between such pair. This process enables the KG to link synonyms across different passages, facilitating the integration of both old and new knowledge during learning. 3) Finally, this phrase-based KG is combined with the original passages, allowing the final open KG to incorporate both conceptual and contextual information (§3.2).
Online Retrieval. 1) The query is linked to relevant triples and passages using the encoder, identifying potential seed nodes for graph search (§3.3). 2) During triple linkage, the recognition memory functions as a filter, ensuring only relevant triples are retained from the retrieved set (§3.4). 3) Given seed nodes, the PPR algorithm is then applied for context-aware retrieval, refining the linking results to retrieve the most relevant passages. 4) Finally, the retrieved passages serve as contextual inputs for the final QA task. Next, we describe each of the improvements in HippoRAG 2 in more detail.
3.2 Dense-Sparse Integration
@@
3.3 Deeper Contextualization
3.4 Recognition Memory
3.5 Online Retrieval
We summarize the online retrieval process in HippoRAG 2 after introducing the above improvements. The task involves selecting seed nodes and assigning reset probabilities for retrieval. HippoRAG 2 identifies phrase nodes from filtered triples generated by query-to-triple and recognition memory. If no triples are available, it directly retrieves top-ranked passages using the embedding model. Otherwise, up to phrase nodes are selected based on their average ranking scores across filtered triples they originate. All passage nodes are also taken as seed nodes, as broader activation improves multi-hop reasoning. Reset probabilities are assigned based on ranking scores for phrase nodes, while passage nodes receive scores proportional to their embedding similarity, adjusted by a weight factor (§6.2) to balance the influence between phrase nodes and passage nodes. The PPR search is then executed, and passages are ranked by their PageRank scores, with the top-ranked passages used for downstream QA. An example of the pipeline is in Appendix B and the PPR initialization is detailed in Appendix G.1,
4 Experimental Setup
4.1 Baselines
We select three types of comparison methods: 1) The classic retrievers BM25 (Robertson & Walker, 1994), Contriever (Izacard et al., 2022) and GTR (Ni et al., 2022). 2) Large embedding models that perform well on the BEIR leaderboard (Thakur et al., 2021), including Alibaba-NLP/GTE-Qwen2-7B-Instruct (Li et al., 2023), GritLM/GritLM-7B (Muennighoff et al., 2024), and nvidia/NV-Embed-v2 (Lee et al., 2025). 3) Structure-augmented RAG methods, including RAPTOR (Sarthi et al., 2024), GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2024), and HippoRAG (Gutiérrez et al., 2024).
4.2 Datasets
To evaluate how well RAG systems retain factual memory while enhancing associativity and sense-making, we select datasets that correspond to three critical challenge types: 1) Simple QA primarily evaluates the ability to recall and retrieve factual knowledge accurately. 2) Multi-hop QA measures associativity by requiring the model to connect multiple pieces of information to derive an answer. 3) Discourse understanding evaluates sense-making by testing the capability to interpret and reason over lengthy, complex narratives. The statistics for our sampled dataset are summarized in Table 1.
@@ 4.2 Datasets
4.3 Metrics
4.4 Implementation Details
For HippoRAG 2, we use the open-source Llama-3.3-70B-Instruct (AI@Meta, 2024) as both the extraction (NER and OpenIE) and triple filtering model, and we use nvidia/NV-Embed-v2 as the retriever. We also reproduce the compared structure-augmented RAG methods using the same extractor and retriever for a fair comparison. For the triple filter, we use DSPy (Khattab et al., 2024) MIPROv2 optimizer and Llama-3.3-70B-Instruct to tune the prompt, including the instructions and demonstrations. The resulting prompt is shown in Appendix A. We use top-5 triples ranked by retriever for filtering. For hyperparameters, we follow the default settings from HippoRAG. More implementation and hyperparameter details can be found in Appendix G.
| Retriever | Dense Retrieval | HippoRAG 2 |
| GTE-Qwen2-7B-Instruct | ||
| GritLM-7B | ||
| NV-Embed-v2 (7B) |
@@ 1. Erik Hort
5 Results
QA Performance. Table 2 presents the QA performance of various retrievers across multiple RAG benchmarks using Llama-3.3-70B-Instruct as the QA reader. HippoRAG 2 achieves the highest average F1 score, demonstrating robustness across different settings. Large embedding models outperform smaller ones, with NV-Embed-v2 (7B) scoring higher on average than GTR (T5-base). These models also surpass structure-augmented RAG methods with lower computational costs but excel mainly in simple QA while struggling in complex cases. N otably, HippoRAG 2 outperforms NV-Embed-v2 by F1 on 2Wiki and by on the challenging LV-Eval dataset. Compared to HippoRAG, HippoRAG 2 shows even greater improvements, validating its neuropsychology-inspired approach. These results highlight HippoRAG 2 as a state-of-the-art RAG system that enhances both retrieval and QA performance while being effectively powered by an open-source model. Table 8 in Appendix C presents additional QA results (EM and F1) using Llama or GPT-4o-mini as the QA reader, along with an extractor or triple filter. GPT-4o-mini follows Llama’s trend, with NV-Embed-v2 outperforming structure-augmented methods in most cases, except for HippoRAG in multi-hop QA. HippoRAG 2 consistently outperforms all other methods across nearly all settings.
Retrieval Performance. We report retrieval results for datasets with supporting passage annotations and models that explicitly retrieve passages in Table 3. Large embedding models (7B) significantly outperform classic smaller LM-based models like Contriever and GTR, achieving at least a higher F1 score. While our reproduction of HippoRAG using Llama-3.3-70B-Instruct and NV-Embed-v2 shows slight improvements over the original paper, the gains are minimal, with only a increase in F1. Although HippoRAG excels in entity-centric retrieval, achieving the highest recall@5 on PopQA, it generally lags behind recent dense retrievers and HippoRAG 2. Notably, HippoRAG 2 achieves the highest recall scores across most datasets, with substantial improvements of and in Recall@5 on MuSiQue and 2Wiki, respectively, compared to the strongest dense retriever, NV-Embed-v2. Additionally, the cost and efficiency analysis is presented in Appendix F.
6 Discussions
6.1 Ablation Study
6.2 Controlling Reset Probabilities
@@ 6.3 Dense Retriever Flexibility
The dense retriever employed by HippoRAG 2 is fully plug-and-play, offering seamless integration. As demonstrated in Table 6, HippoRAG 2 consistently surpasses direct dense retrieval across various retrievers. Notably, these performance gains remain robust regardless of the specific dense retriever used.
@@ Appendix F Cost and Efficiency
For LLM deployment, we run Llama-3.3-70B-Instruct on a machine equipped with four NVIDIA H100 GPUs, utilizing tensor parallelism via vLLM (Kwon et al., 2023). We also employ the gpt-4o-mini-2024-07-18 model from OpenAI’s official endpoint, leveraging its batch API222https://platform.openai.com/docs/guides/batch. For offline indexing, we execute NER and Open IE on the MuSiQue corpus ( passages). Processing each passage takes approximately seconds using Llama-3.3-70B-Instruct, while utilizing the gpt-4o-mini batch API allows indexing to complete within 24 hours at a cost of under $ USD.
Comparison With Structure-Augmented RAG Methods
We count the token usage across different structure-augmented RAG methods when indexing the MuSiQue corpus using the Llama-3.3-70B-Instruct model, and we compare the number of input and output tokens against RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024), and GraphRAG (Edge et al., 2024) in Table 12. HippoRAG 2 not only outperforms these RAG methods in QA and retrieval performance but also uses much fewer tokens compared to LightRAG and GraphRAG.
| HippoRAG 2 | RAPTOR | LightRAG | GraphRAG | |
| Input Tokens | ||||
| Output Tokens |

### v2

@@ Abstract
Our ability to continuously acquire, organize, and leverage knowledge is a key feature of human intelligence that AI systems must approximate to unlock their full potential. Given the challenges in continual learning with large language models (LLMs), retrieval-augmented generation (RAG) has become the dominant way to introduce new information. However, its reliance on vector retrieval hinders its ability to mimic the dynamic and interconnected nature of human long-term memory. Recent RAG approaches augment vector embeddings with various structures like knowledge graphs to address some of these gaps, namely sense-making and associativity. However, their performance on more basic factual memory tasks drops considerably below standard RAG. We address this unintended deterioration and propose HippoRAG 2, a framework that outperforms standard RAG comprehensively on factual, sense-making, and associative memory tasks. HippoRAG 2 builds upon the Personalized PageRank algorithm used in HippoRAG and enhances it with deeper passage integration and more effective online use of an LLM. This combination pushes this RAG system closer to the effectiveness of human long-term memory, achieving a 7% improvement in associative memory tasks over the state-of-the-art embedding model while also exhibiting superior factual knowledge and sense-making memory capabilities. This work paves the way for non-parametric continual learning for LLMs. Code and data are available at https://github.com/OSU-NLP-Group/HippoRAG.
@@ 2.1 Continual Learning for LLMs
As the use of LLMs in real-world applications grows, it becomes increasingly important for them to acquire and integrate new knowledge over time while preserving past information—as evidenced by the many benchmarking efforts in this direction (Zhong et al., 2023; Liska et al., 2022; Kim et al., 2024; Roth et al., 2024; Li et al., 2024). Given the high computational cost of full-scale LLM pretraining, various techniques have been leveraged to endow these models with this continual learning capacity. These approaches generally fall into three categories: continual fine-tuning, model editing, and RAG (Shi et al., 2024).
@@ 2.2 Non-Parametric Continual Learning for LLMs
Encoder model improvements, particularly with LLM backbones, have significantly enhanced RAG systems by generating high-quality embeddings that better capture semantic relationships, improving retrieval quality for LLM generation. Recent models (Li et al., 2023; Muennighoff et al., 2025; Lee et al., 2025) leverage LLMs, large corpora, improved architectures, and instruction fine-tuning for notable retrieval gains. NV-Embed-v2 (Lee et al., 2025) serves as the primary comparison in this paper.
@@ 3 HippoRAG 2
3.1 Overview
HippoRAG (Gutiérrez et al., 2024) is a neurobiologically inspired long-term memory framework for LLMs in which each component is inspired by its neurobiological analog for human memory. The framework consists of three primary components: 1) an LLM that acts as an artificial neocortex, 2) a KG and the Personalized PageRank algorithm to mirror the auto-associative qualities of the hippocampus and 3) a retrieval encoder that links these two components, reflecting one of the functions of the parahippocampal regions. These components collaborate to replicate the interactions observed in human long-term memory.
HippoRAG’s offline indexing process uses an LLM to process passages into KG triples, which are then incorporated into the KG, our artificial hippocampal index. Meanwhile, the retrieval encoder is responsible for detecting synonymy to interconnect information. In HippoRAG’s online retrieval process, the LLM neocortex extracts named entities from a query while the retrieval encoder finds their most similar counterparts in the KG. Then, the nodes in the KG corresponding to these entities, which we refer to as seed nodes, are used to run the Personalized PageRank (PPR) algorithm. More specifically, these seed nodes are used to assign the reset probabilities within PPR, which alter the original PageRank algorithm to distribute probability towards the seed nodes and their neighborhoods, enabling HippoRAG’s context-based retrieval. Although HippoRAG seeks to construct memory from non-parametric RAG, its effectiveness is hindered by a critical flaw: an entity-centric approach that causes context loss during both indexing and inference, as well as difficulties in semantic matching.
Built on the neurobiologically inspired long-term memory framework proposed in HippoRAG (Gutiérrez et al., 2024), the structure of HippoRAG 2 follows a similar two-stage process: offline indexing and online retrieval, as shown in Figure 2. Additionally, however, HippoRAG 2 introduces several key refinements that improve its alignment with human memory mechanisms: 1) It seamlessly integrates conceptual and contextual information within the KG, enhancing the comprehensiveness and atomicity of the constructed index (§3.2). 2) It facilitates more context-aware retrieval by leveraging the KG structure beyond isolated KG nodes (§3.3). 3) It incorporates recognition memory to improve seed node selection for graph search (§3.4). In the following sections, we introduce the pipeline in more detail and elaborate on each of these refinements.
Offline Indexing. 1) HippoRAG 2, just as HippoRAG, leverages an LLM to extract triples from each passage using OpenIE, which allows the relations and entities to be generated without any constraints or schema. These triples are then arranged into our schema-less KG or hippocampal index. We call the subject or object of these triples phrases and the edge connecting them relation edge. 2) Next, the retrieval encoder identifies synonyms by evaluating phrase pairs within the KG, detecting those with vector similarity above a predefined threshold, and adding synonym edge between such pair. This process enables the KG to link synonyms across different passages, facilitating the integration of both old and new knowledge during learning. 3) Finally, this phrase-based KG is combined with the original passages, allowing the final open KG to incorporate both conceptual and contextual information (§3.2).
Online Retrieval. 1) The query is linked to relevant triples and passages using the encoder, identifying nodes that could be used as seed nodes for graph search (§3.3). 2) During triple linkage, the recognition memory functions as a filter, ensuring only relevant triples are retained from the retrieved set as the final seed nodes (§3.4). 3) These final seed nodes are then used to assign reset probabilities within the PPR algorithm, enabling its context-aware retrieval and refining the linking results to retrieve the most relevant passages. 4) Finally, the retrieved passages serve as contextual inputs for the final QA task. Next, we describe each of the improvements in HippoRAG 2 in more detail.
3.2 Dense-Sparse Integration
3.3 Deeper Contextualization
3.4 Recognition Memory
3.5 Online Retrieval
We summarize the online retrieval process in HippoRAG 2 after introducing the above improvements. The task involves selecting seed nodes and assigning reset probabilities for retrieval. HippoRAG 2 identifies phrase nodes from filtered triples generated by query-to-triple and recognition memory. If no triples are available, it directly retrieves top-ranked passages using the embedding model. Otherwise, up to phrase nodes are selected based on their average ranking scores across filtered triples they originate. All passage nodes are also taken as seed nodes, as broader activation improves multi-hop reasoning. Reset probabilities are assigned based on ranking scores for phrase nodes, while passage nodes receive scores proportional to their embedding similarity, adjusted by a weight factor (§6.2) to balance the influence between phrase nodes and passage nodes. The PPR search is then executed, and passages are ranked by their PageRank scores, with the top-ranked passages used for downstream QA. An example of the pipeline is in Appendix B and the PPR initialization is detailed in Appendix G.1.
@@ 3.5 Online Retrieval
4 Experimental Setup
4.1 Baselines
We select three different types of baselines for comparison. We include three simple baselines: the classic BM25 (Robertson & Walker, 1994) baseline as well as Contriever (Izacard et al., 2022) and GTR (Ni et al., 2022), two popular dense embedding retrievers.
Our second baseline category includes some of the largest embedding models available (7B) that demonstrate strong performance on the BEIR leaderboard (Thakur et al., 2021): Alibaba-NLP/GTE-Qwen2-7B-Instruct (Li et al., 2023), GritLM/GritLM-7B (Muennighoff et al., 2025), and nvidia/NV-Embed-v2 (Lee et al., 2025).
In our final baseline category, we include four structure-augmented RAG methods. RAPTOR (Sarthi et al., 2024) organizes the retrieval corpus into a hierarchical structure based on semantic similarity. GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024) leverage a KG structure like ours to generate high-level summaries of the concepts present in the corpus. Finally, HippoRAG (Gutiérrez et al., 2024) uses a KG as well but integrates knowledge using PPR rather than summarization.
4.2 Datasets
To evaluate how well RAG systems retain factual memory while enhancing associativity and sense-making, we select datasets that correspond to three critical challenge types.
1.
Simple QA primarily evaluates the ability to recall and retrieve factual knowledge accurately.
2.
Multi-hop QA measures associativity by requiring the model to connect multiple pieces of information to derive an answer.
3.
Discourse understanding evaluates sense-making by testing the capability to interpret and reason over lengthy, complex narratives.
We will now list the datasets chosen for each category and describe them in detail. The statistics for our sampled datasets are summarized in Table 1.
@@ 4.2 Datasets
4.3 Metrics
4.4 Implementation Details
For HippoRAG 2, we use the open-source Llama-3.3-70B-Instruct (AI@Meta, 2024) as both the extraction (NER and OpenIE) and triple filtering model, and we use nvidia/NV-Embed-v2 as the retriever. We also reproduce the compared structure-augmented RAG methods using the same extractor and retriever for a fair comparison. For the triple filter, we use DSPy (Khattab et al., 2024) MIPROv2 optimizer and Llama-3.3-70B-Instruct to tune the prompt, including the instructions and demonstrations. The resulting prompt is shown in Appendix A. We use top-5 triples ranked by retriever for filtering. Our QA module uses the top-5 retrieved passages as context for an LLM (GPT-4o-mini or Llama-3.3-70B-Instruct) to generate the final answer. For hyperparameters, we follow the default settings from HippoRAG. More implementation and hyperparameter details can be found in Appendix G.
5 Results
QA Performance. Table 2 presents the QA performance of various retrievers across multiple RAG benchmarks using Llama-3.3-70B-Instruct as the QA reader. HippoRAG 2 achieves the highest average F1 score, demonstrating robustness across different settings. Large embedding models outperform smaller ones, with NV-Embed-v2 (7B) scoring higher on average than GTR (T5-base). These models also surpass structure-augmented RAG methods with lower computational costs but excel mainly in simple QA while struggling in complex cases. Notably, HippoRAG 2 outperforms NV-Embed-v2 by F1 on 2Wiki and by on the challenging LV-Eval dataset. Compared to HippoRAG, HippoRAG 2 shows even greater improvements, validating its neuropsychology-inspired approach. These results highlight HippoRAG 2 as a state-of-the-art RAG system that enhances both retrieval and QA performance while being effectively powered by an open-source model. Table 8 in Appendix C presents additional QA results (EM and F1) using Llama or GPT-4o-mini as the QA reader, along with an extractor or triple filter. GPT-4o-mini follows Llama’s trend, with NV-Embed-v2 outperforming structure-augmented methods in most cases, except for HippoRAG in multi-hop QA. HippoRAG 2 consistently outperforms all other methods across nearly all settings. An analysis of the computational resources (tokens, time and memory) required for each method can be found in Appendix F.
Retrieval Performance. We report retrieval results for datasets with supporting passage annotations and models that explicitly retrieve passages in Table 3. Large embedding models (7B) significantly outperform classic smaller LM-based models like Contriever and GTR, achieving at least a higher F1 score. While our reproduction of HippoRAG using Llama-3.3-70B-Instruct and NV-Embed-v2 shows slight improvements over the original paper, the gains are minimal, with only a increase in F1. Although HippoRAG excels in entity-centric retrieval, achieving the highest recall@5 on PopQA, it generally lags behind recent dense retrievers and HippoRAG 2. Notably, HippoRAG 2 achieves the highest recall scores across most datasets, with substantial improvements of and in Recall@5 on MuSiQue and 2Wiki, respectively, compared to the strongest dense retriever, NV-Embed-v2.
@@ 5 Results
6 Discussions
6.1 Ablation Study
6.2 Controlling Reset Probabilities
6.3 Robustness to Corpus Expansion
As RAG systems become more widely adopted in the real-world, they must increasingly adapt to continual learning scenarios in which the retrieval corpora grow continuously. To understand how HippoRAG 2’s capacity to handle this setting compared to standard RAG, we design an experiment in which we partition NQ and MuSiQue into four equal segments, each containing the gold documents and distractors for approximately 250 questions. We then select one segment for evaluation and incrementally add the remaining segments, measuring how performance evolves as new knowledge is added, allowing us to simulate a continual learning setting. We show the F1 score for HippoRAG 2 and NV-Embed-v2, our strongest baseline, in Figure 3.
@@ 1. Erik Hort
| Retriever | Dense Retrieval | HippoRAG 2 |
| GTE-Qwen2-7B-Instruct | ||
| GritLM-7B | ||
| NV-Embed-v2 (7B) |
As we can see in Figure 3, HippoRAG 2’s improvements over NV-Embed-v2 remain remarkably consistent in both simple (NQ) and associative (MuSiQue) continual learning settings. We also note that, while both methods retain strong performance on simple QA (solid lines) as more knowledge is introduced, their performance in the more complex associative task (dotted lines) degrades at a similar rate as more information is introduced. This divergence underscores the importance of incorporating varied task complexities into future continual learning benchmarks.
@@ 6.4 Dense Retriever Flexibility
As demonstrated in Table 7, HippoRAG 2 consistently surpasses direct dense retrieval across various retrievers. Notably, these performance gains remain robust regardless of the specific dense retriever used.
@@ Appendix F Cost and Efficiency
For LLM deployment, we run Llama-3.3-70B-Instruct on a machine equipped with four NVIDIA H100 GPUs, utilizing tensor parallelism via vLLM (Kwon et al., 2023).
For a detailed comparison with our baselines, we track the computational resources (# of tokens, indexing time, time per query and GPU memory) usage when indexing and performing QA on the MuSiQue corpus (11k documents) using the Llama-3.3-70B-Instruct model. We compare HippoRAG 2 against NV-Embed-v2 (Lee et al., 2025), RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024), HippoRAG (Gutiérrez et al., 2024) and GraphRAG (Edge et al., 2024) in Table 12. For the memory requirements, we ignore all memory for model weights since it is shared across all systems.
HippoRAG 2 not only outperforms these RAG methods in QA and retrieval performance but also uses much fewer tokens compared to LightRAG and GraphRAG. In terms of time, we note that HippoRAG 2 is much more efficient than GraphRAG and LightRAG while only being slightly less efficient than both RAPTOR and HippoRAG. HippoRAG 2 ’s use of fact embeddings does increase its memory requirements compared to our baselines, however, we believe that this is an acceptable tradeoff given our method’s performance benefits. Additionally, while all approaches lag behind standard RAG in terms of time and memory efficiency, HippoRAG 2 is the only one that outperforms this strong baseline substantially.
| NV-Embed-v2 | RAPTOR | LightRAG | GraphRAG | HippoRAG | HippoRAG 2 | |
| Input Tokens | – | |||||
| Output Tokens | – | |||||
| Indexing Time (min) | ||||||
| QA Time/Query (sec) | ||||||
| QA GPU Memory (GB) |